OpenStudy (anonymous):

To which subsets of real numbers does the number -22 belong? Choose all subsets that apply. theres only 2 btw A.Whole Numbers B.Rational Numbers C.Integers D.Irrational Numbers E.Natural Numbers

OpenStudy (mathstudent55):

Natural Numbers (also called Counting Numbers): 1, 2, 3, 4, 5, ... Whole Numbers: 0, 1, 2, 3, 4, 5, ... Integers: ..., -3, -2, -1, 0, 1, 2, 3, ... Rational Number: a number that can be written as a fraction of integers Irrational Number: a number that is not rational

OpenStudy (anonymous):

hmm

OpenStudy (mathstudent55):

Go through each definition, and see if -22 belongs in each set. Notice that the definitions are not in the same order as your choices.

OpenStudy (anonymous):

okay so the first is integers but i dont understand what the second one would be

OpenStudy (anonymous):

?

OpenStudy (anonymous):

the first is integers but i dont know the second one

OpenStudy (mathstudent55):

-22 is definitely an integer, but it is part of another set in the choices too.

OpenStudy (anonymous):

irrational numbers?

OpenStudy (mathstudent55):

No. Look at the definition of rational. Can -22 be written as a fraction with an integer in the numerator and an integer in the denominatior?

OpenStudy (mathstudent55):

For example, \(\dfrac{-22}{1} \) Is this fraction equal to -22? Are -22 and 1 integers?

OpenStudy (anonymous):

Hmm so is it rational?

OpenStudy (mathstudent55):

Correct. -22 can be written as a fraction of integers in may ways. Here is another way: \(\dfrac{-44}{2} \) -44 and 2 are both integers, and that fraction equals -22, so -22 is a rational number.

OpenStudy (anonymous):

Awesome :D

OpenStudy (anonymous):

Is it okay if i copy and paste ur notes to my notebook? for learning

OpenStudy (mathstudent55):

Yes, be my guest. Your problem is already answered, but for a little more understanding, read on. Now we can ask a question. Then what is an irrational number? A number such as \(\pi\) which is the circumference divided by the diameter of a circle. \(\sqrt 2\), \(\sqrt 3\) These numbers cannot be written as a fraction of integers. You may have seen an approximation of \(\pi\) written as \(\dfrac{22}{7} \), but that is only a rational approximation. \(\pi\) is an irrational number.

OpenStudy (anonymous):

Okay well thank you for your help :)

OpenStudy (mathstudent55):

You're welcome.